# The Diophantine problem in isotropic reductive groups

### v1

@@ 1 Introduction
$\exists\vec{y}\enskip\bigwedge_{i=1}^{n}P_{i}(\vec{x},\vec{y})$
@@
A closely related is the notion of e-interpretation. An e-interpretation of a model $M$ in a model $N$ (of possibly another language) is a surjection $X\to M$ for some Diophantine $X\subseteq N^{n}$ such that the lifts of all relations on $M$ to $X$ from its language are Diophantine (including the kernel pair lifting the equality), as well as lifts of all graphs of operations on $M$. For example, if $K$ is a commutative ring and $G$ is a finitely presented affine group scheme over $K$, then the group $G(K)$ is e-interpretable in the ring $K$.
Suppose that $M$ and $L$ are enumerated by $\mathbb{N}$ or its finite subsets. The Diophantine problem $\mathcal{D}(M)$ is decidable if the non-emptiness of generic Diophantine set $\{\vec{x}\mid\varphi(\vec{x},\vec{a})\}$ may be checked by an algorithm knowing only the codes of $\varphi$ and $\vec{a}$ (note that the set of Diophantine formulae is decidable). For example, the Diophantine problem $\mathcal{D}(\mathbb{Z})$ (for the language of rings and a natural enumeration) is undecidable by famous Matiyasevich’s theorem. On the other hand, $\mathcal{D}(\mathbb{Q}^{\mathrm{alg}})$ is decidable (again for the language of rings and a natural enumeration) since every algebraically closed field admits quantifier elimination.
Now let $K$ be a commutative ring and $G$ be a reductive group scheme over $K$ in the sense of [3]. If $G$ is splits, i.e. it is a Chevalley–Demazure group scheme, simple, and of rank at least $2$, then by the main result of Elena Bunina, Alexey Myasnikov, and Eugene Plotkin [2] the Diophantine problems $\mathcal{D}(K)$ and $\mathcal{D}(G(K))$ are equivalent. More precisely, if $K$ is countable with fixed enumeration, then these two problems are reducible to each other by explicit algorithms.
@@
We follow the general strategy of Bunina – Myasnikov – Plotkin. Namely, root subgroups of $G(K)$ turn out to be Diophantine, see theorem 4 below. This allows us to construct e-interpretation of the base ring $K$ in the group $G(K)$, this is theorem 5. Finally, since $K$ and $G(K)$ e-interpret each other, their Diophantine problems are equivalent (theorem 6). As an application of preparatory technical results we also find the centralizer and the normalizer of all root subgroups together, this is a generalization of [1] to the isotropic case.
@@ 2 Isotropic reductive groups
Let $K$ be a commutative ring (all rings in this paper are unital) and $G$ a reductive group scheme over $K$ [3, Exp. XIX]. By root systems we mean crystallographic root systems possibly of type $\mathsf{BC}_{\ell}$. Recall that there is an étale extension $K\subseteq\widetilde{K}$ such that $G_{\widetilde{K}}$ splits with possibly non-constant root datum, i.e. there is a decomposition $\widetilde{K}=\widetilde{K}_{1}\times\ldots\times\widetilde{K}_{n}$ such that each $G_{\widetilde{K}_{i}}$ is isomorphic to a split reductive group scheme $\mathrm{G}(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi};\widetilde{K}_{i})$ over $\widetilde{K}_{i}$ [3, corollary XXII.2.3]. We always assume that $n=1$, i.e. the root datum $(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi})$ is constant over $\widetilde{K}$, and that the root system $\widetilde{\Phi}$ is irreducible. But $G$ is not necessarily semisimple, i.e. the scheme center $\operatorname{C}(G)$ may contain a non-trivial torus (such generality is required e.g. in applications of lemma 5 below).
In this paper we call $G$ isotropic with root system $\Phi$ if there are subgroups $L\leq G$ and $G_{\alpha}\leq G$ for $\alpha\in\Phi$ such that after some (each sufficiently large) étale extension $K\subseteq\widetilde{K}$ there are an isomorphism
$F\colon G_{\widetilde{K}}\to\mathrm{G}(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi};\widetilde{K})$
with a split reductive group scheme and a map $u\colon\widetilde{\Phi}\to\Phi\sqcup\{0\}$ induced by a linear map of the ambient vector spaces such that
  • $\Phi$ is contained in the image of $u$;
  • $F(G_{\alpha,\widetilde{K}})=\langle U_{\beta}\mid u(\beta)\in\{\alpha,2\alpha\}\rangle$ is a standard unipotent subgroup, where $U_{\beta}$ are the root subgroups of the split reductive group scheme;
  • $F(L_{\widetilde{K}})=\mathrm{G}(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi};\widetilde{K})^{0}_{u^{-1}(0)}$ is a reductive subsystem subgroup, namely, the subgroup of type (RC) corresponding to the subset $u^{-1}(0)$ [3, §XXII.5.11];
  • the group $G$ together with the family $(G_{\alpha})_{\alpha\in\Phi}$ is $\Phi$-graded, i.e. $G_{2\alpha}\leq G_{\alpha}$ for ultrashort $\alpha$,
    $[G_{\alpha},G_{\beta}]\leq\langle G_{\gamma}\mid\gamma\in(\mathbb{N}_{+}\alpha+\mathbb{N}_{+}\beta)\cap\Phi\rangle$
    (this automatically follows from the previous condition), and there are Weyl elements
    $n_{\alpha}\in G_{\alpha}(K)\,G_{-\alpha}(K)\,G_{\alpha}(K)$
    for all $\alpha\in\Phi$, i.e. elements with the property ${{}^{n_{\alpha}}\!{G_{\beta}}}=G_{s_{\alpha}(\beta)}$;
  • The map $u\colon\widetilde{\Phi}\to\Phi\sqcup\{0\}$ comes from one of irreducible Tits indices.
For example, all reductive groups over semi-local rings (and even LG-rings [4]) with connected spectra are naturally isotropic using their minimal parabolic subgroups by [3, Exp. XXVI] and [6]. All these isotropic structures are conjugate by elements of $G(K)$.
@@
Recall [7] that an irreducible Tits index $(\widetilde{\Phi},\Gamma,J)$ consists of a reduced irreducible root system $\widetilde{\Phi}$, a subgroup $\Gamma$ of its group of outer automorphisms (i.e. the automorphism group of its Dynkin diagram), and a $\Gamma$-invariant subset $J$ of vertices of the Dynkin diagram satisfying an additional condition (namely, that it may be constructed by a reductive group scheme over a field using its minimal parabolic subgroup). Then $\Phi$ is the image of $\widetilde{\Phi}$ in the factor-space of the ambient vector space by $\Gamma$ and the span of basic roots not in $J$, the rank of $\Phi$ is the number of $\Gamma$-orbits in $J$. An index is usually denoted as ${{}^{g}\!{\mathsf{X}_{n,r}^{t}}}$, where $g=|\Gamma|$, $\mathsf{X}_{n}$ is the type of $\widetilde{\Phi}$, $r$ is the rank of $\Phi$, and $t$ is an additional parameter (see the list below).
Here is the list of irreducible Tits indices using the standard (Bourbaki) numeration of simple roots.
  • ${{}^{1}\!{\mathsf{A}_{n,r}^{(d)}}}$, where $d(r+1)=n+1$ and $n\geq 1$. Here $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{A}_{r}$, modulo the center $G(K)=\mathrm{PGL}(r+1,A)$ for an Azumaya algebra $A$ over $K$ of rank $d$. Also, $u^{-1}(0)=(r+1)A_{d-1}$ and $|u^{-1}(\alpha)|=d^{2}$ for every root $\alpha$.
  • ${{}^{2}\!{\mathsf{A}_{n,r}^{(d)}}}$, where $d\mid n+1$, $n\geq 1$, and $12rd\leq n+$ 1. Here
    $J=\{d,2d,\ldots,rd;n+1-rd,\ldots,n+1-2d,n+1-d\}.$
    The relative root system is $\Phi=\mathsf{C}_{r}$ for $12rd=n+$ 1 and $\Phi=\mathsf{BC}_{r}$ otherwise. Modulo the center $G(K)=\mathrm{SU}(A,h)$ for an Azumaya algebra $A$ of rank $d$ over a quadratic étale extension $K\subseteq K^{\prime}$ and for a non-degenerate hermitian form $h$ over $A$ of rank $\frac{n+1}{d}$ and Witt index $r$ (i.e. with a chosen hyperbolic subspace of rank $2r$). Also, $u^{-1}(0)=2r\mathsf{A}_{d-1}+\mathsf{A}_{n-2rd}$ (the last summand is vacuous for $2rd\in\{n,n+1\}$), $|u^{-1}(\alpha)|=d^{2}$ for long $\alpha$, $|u^{-1}(\alpha)|=2d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n+1-2rd)$ for ultrashort $\alpha$ (and $2rd\leq n$).
  • $\mathsf{B}_{n,r}$, where $n\geq 2$ and $r\leq n$. Here $J=\{1,2,\ldots,r\}$, $\Phi=\mathsf{B}_{r}$, and $G(K)=\mathrm{SO}(K,q)$ modulo the center, where $q$ is a semi-regular quadratic form of rank $12n+$ 1 and Witt index $r$. Also, $u^{-1}(0)=\mathsf{B}_{n-r}$, $|u^{-1}(\alpha)|=1$ for long $\alpha$, and $|u^{-1}(\alpha)|=2n+1-2r$ for short $\alpha$.
  • $\mathsf{C}_{n,r}^{(d)}$, where $n\geq 3$, $d=2^{k}\mid 2n$, $rd\leq n$, and $n=r$ in the case $d=1$. Here $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{C}_{r}$ for $rd=n$ and $\Phi=\mathsf{BC}_{r}$ otherwise. Modulo the center $G(K)=\mathrm{U}(A,h)$, where $A$ is an Azumaya algebra over $K$ of rank $d$ with symplectic involution, $h$ is a non-degenerate anti-hermitian form of rank $\frac{2n}{d}$ and Witt index $r$ (with the additional condition that $h$ comes from a symplectic form, not an arbitrary alternating form). Also, $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{C}_{n-rd}$, $|u^{-1}(\alpha)|=\frac{d(d+1)}{2}$ for long $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$ (and $rd<n$).
  • ${{}^{1}\!{\mathsf{D}_{n,r}^{(d)}}}$, where $n\geq 4$, $d=2^{k}\mid 2n$, $rd\leq n$, and $n\neq rd+1$. Here $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{D}_{r}$ for $rd=n$ and $d=1$, $\Phi=\mathsf{C}_{r}$ for $rd=n$ and $d>1$, $\Phi=\mathsf{B}_{r}$ for $rd<n$ and $d=1$, and $\Phi=\mathsf{BC}_{r}$ for $rd<n$ and $d>1$. Modulo the center $G(K)=\mathrm{SU}(A,h,q)$, where $A$ is an Azumaya algebra over $K$ of rank $d$ with orthogonal involution, $h$ is a non-degenerate hermitian form of rank $\frac{2n}{d}$ and Witt index $r$, and $q$ is a suitable associated quadratic form. Also, $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{D}_{n-rd}$, $|u^{-1}(\alpha)|=\frac{d(d-1)}{2}$ for long $\alpha$ (and $d>1$), $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$ (and $rd<n$). Here roots of the relative root systems of types $\mathsf{D}_{r}$ and $\mathsf{B}_{r}$ are called short and ultrashort depending on their length, not long and short.
  • ${{}^{2}\!{\mathsf{D}_{n,r}^{(d)}}}$, where $n\geq 4$, $d=2^{k}\mid 2n$, and $rd\leq n-1$. Here $J=\{d,2d,\ldots,rd\}$ for $rd<n-1$ and $J=\{d,2d,\ldots,n-1,n\}$ otherwise, $\Phi=\mathsf{BC}_{r}$ for $d>1$ and $\Phi=\mathsf{B}_{r}$ otherwise. Modulo the center $G(K)=\mathrm{SU}(A,h,q)$, where $A$ is an Azumaya algebra over $K$ of rank $d$ with orthogonal involution, $h$ is a non-degenerate hermitian form of rank $\frac{2n}{d}$ and Witt index $r$, and $q$ is a suitable associated quadratic form. Also, $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{D}_{n-rd}$, $|u^{-1}(\alpha)|=\frac{d(d-1)}{2}$ for long $\alpha$ (and $d>1$), $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$. As in the previous case, roots of the relative root system of type $\mathsf{B}_{r}$ are called short and ultrashort.

### v2

@@ 1 Introduction
$\exists\vec{y}\enskip\bigl(P_{1}(\vec{x},\vec{y})\wedge\ldots\wedge P_{n}(\vec{x},\vec{y})\bigr)$
@@
A closely related is the notion of e-interpretation. An e-interpretation of a model $M$ in a model $N$ (of possibly another language) is a surjection $X\to M$ for some Diophantine $X\subseteq N^{n}$ such that the lifts of all relations on $M$ to $X$ from its language are Diophantine (including the equivalence relation lifting the equality), as well as lifts of all graphs of operations on $M$. For example, if $K$ is a commutative ring and $G$ is a finitely presented affine group scheme over $K$, then the group $G(K)$ is e-interpretable in the ring $K$.
Suppose that $M$ and $L$ are enumerated by $\mathbb{N}$ or its finite subsets. The Diophantine problem $\mathcal{D}(M)$ is decidable if the non-emptiness of generic Diophantine set $\{\vec{x}\mid\varphi(\vec{x},\vec{a})\}$ may be checked by an algorithm knowing only the codes of $\varphi$ and $\vec{a}$ (note that the set of Diophantine formulae is decidable). For example, the Diophantine problem $\mathcal{D}(\mathbb{Z})$ (for the language of rings and a natural enumeration) is undecidable by famous Matiyasevich’s theorem. On the other hand, $\mathcal{D}(\mathbb{Q}^{\mathrm{alg}})$ is decidable (again for the language of rings and a natural enumeration) because every algebraically closed field admits quantifier elimination.
Now let $K$ be a commutative ring and $G$ be a reductive group scheme over $K$ in the sense of [13]. If $G$ is splits, i.e. it is a Chevalley–Demazure group scheme, simple, and of rank at least $2$, then by the main result of Elena Bunina, Alexey Myasnikov, and Eugene Plotkin [12] the ring $K$ and the group $G(K)$ are e-interpretable in each other and the Diophantine problems $\mathcal{D}(K)$ and $\mathcal{D}(G(K))$ are equivalent. More precisely, if $K$ is countable with fixed enumeration, then the Diophantine problems are reducible to each other by explicit algorithms. See also [2, 3, 4, 5, 7, 8, 9, 10, 11, 14, 16, 17, 18, 19, 23] for other results on model theory of linear groups over rings, including e-interpretability and first order interpretability.
@@
We follow the general strategy of Bunina – Myasnikov – Plotkin. Namely, root subgroups of $G(K)$ turn out to be Diophantine, see theorem 4 below. This allows us to construct e-interpretation of the base ring $K$ in the group $G(K)$, this is theorem 5. Finally, since $K$ and $G(K)$ e-interpret each other, their Diophantine problems are equivalent (theorem 6). As an application of preparatory technical results we also find the centralizer and the normalizer of all root subgroups (and even their finite subsets) together, this is a generalization of [1, 15].
@@ 2 Isotropic reductive groups
In this paper all root systems are crystallographic, but not necessarily reduced, i.e. we allow components of type $\mathsf{BC}_{\ell}$. Let $K$ be a non-zero unital commutative ring and $\widetilde{\Phi}$ be a reduced root system. Choose a weight lattice
$\mathbb{Z}\widetilde{\Phi}\leq\Lambda\leq\bigl\{x\mid\forall\alpha\in\widetilde{\Phi}\enskip 2\frac{x\cdot\alpha}{\alpha\cdot\alpha}\in\mathbb{Z}\bigr\}$
[…]
$L=\bigl\langle\widetilde{T},\widetilde{U}_{\alpha}\mid u(\alpha)=0\bigr\rangle,\quad U_{\alpha}=\bigl\langle\widetilde{U}_{\alpha^{\prime}}\mid u(\alpha^{\prime})\in\{\alpha,2\alpha\}\bigr\rangle$
[…]
This class of group schemes is interesting because every simple reductive group scheme $G$ over a local ring $K$ is anisotropic or isotropic in the above sense in a canonical way up to conjugation, where
$L\rtimes\bigl\langle U_{\alpha}\mid\alpha\in\Phi^{+}\bigr\rangle$
is a minimal parabolic subgroup, see [13, XXVI] and [21] for details. Moreover, if $G$ is a locally isotropic simple reductive group scheme over arbitrary ring $K$ and $\mathfrak{p}\trianglelefteq K$ is a prime ideal, then the isotropic structure of $G_{\mathfrak{p}}$ is well defined on $G_{s}$ for sufficiently large $s\in K\setminus\mathfrak{p}$.
@@
Recall [24] that an irreducible Tits index $(\widetilde{\Phi},\Gamma,J)$ consists of a reduced irreducible root system $\widetilde{\Phi}$, a subgroup $\Gamma$ of its group of outer automorphisms (i.e. the automorphism group of its Dynkin diagram), and a $\Gamma$-invariant subset $J$ of vertices of the Dynkin diagram from the explicit list [24, table II] or [21, appendix]. Then $\Phi$ is the image of $\widetilde{\Phi}$ in the factor-space of the ambient vector space by $\Gamma$ and the span of basic roots not in $J$, the rank of $\Phi$ is the number of $\Gamma$-orbits in $J$. An index is usually denoted as $^{g}\!{\mathsf{X}_{n,r}^{t}}$}, where $g=|\Gamma|$, $\mathsf{X}_{n}$ is the type of $\widetilde{\Phi}$, $r$ is the rank of $\Phi$, and $t$ is an additional parameter. We are interested only in isotropic Tits indices, i.e. with $r\neq 0$. An outer automorphism group of an irreducible Tits index is the group of such automorphisms $f$ of the root system $\widetilde{\Phi}$ that $f$ preserves the basis and $u\circ f=u$.
Now we give the list of all isotropic irreducible Tits indices. It is divided in two parts, classical indices (forming infinite families) and finitely many exceptional ones. The numeration of simple roots in all Dynkin diagrams follows Bourbaki [6]. For classical root systems we use the conventions $\mathsf{A}_{-1}=\mathsf{A}_{0}=\mathsf{B}_{0}=\mathsf{C}_{0}=\mathsf{BC}_{0}=\mathsf{D}_{0}=\mathsf{D}_{1}=\varnothing$ and $\mathsf{D}_{2}=2\mathsf{A}_{1}$, as well as the exceptional isomorphisms $\mathsf{A}_{1}=\mathsf{B}_{1}=\mathsf{C}_{1}$, $\mathsf{B}_{2}=\mathsf{C}_{2}$, $\mathsf{A}_{3}=\mathsf{D}_{3}$.
  • $^{1}\!{\mathsf{A}_{n,r}^{(d)}}$}, where $d(r+1)=n+1$, $d\geq 1$, and $n\geq 1$. Here $\widetilde{\Phi}=\mathsf{A}_{n}$, $J=\{d,2d,\ldots,rd\}$, and $\Phi=\mathsf{A}_{r}$. The kernel of $u$ has the type $(r+1)A_{d-1}$ and $|u^{-1}(\alpha)|=d^{2}$ for every root $\alpha$. The outer automorphism group is trivial for $r\geq 2$ or $^{1}\!{\mathsf{A}_{1,1}^{(1)}}$}, it is cyclic of order $2$ for $r=1$ and $d\geq 2$.
    Such split isotropic structure on $\mathbb{SL}_{n+1}$ corresponds to the block decomposition by $d\times d$ blocks, $U_{\alpha}$ are the groups subschemes of matrices differing from $1$ only in one non-diagonal block, and $L$ is the group subscheme of block diagonal matrices.
  • $^{2}\!{\mathsf{A}_{n,r}^{(d)}}$}, where $d\mid n+1$, $2rd\leq n+1$, $d\geq 1$, $r\geq 1$, and $n\geq 2$. Here $\widetilde{\Phi}=\mathsf{A}_{n}$,
    $J=\{d,2d,\ldots,rd\}\sqcup\{n+1-rd,n+1-(r-1)d,\ldots,n+1-d\}$
    for $2rd\leq n$ and $J=\{d,2d,\ldots,n+1-d\}$ for $2rd=n+1$, $\Phi=\mathsf{BC}_{r}$ for $2rd\leq n$ and $\Phi=\mathsf{C}_{r}$ for $2rd=n+1$. The kernel of $u$ has the type, $2r\mathsf{A}_{d-1}+\mathsf{A}_{n-2rd}$, $|u^{-1}(\alpha)|=2d(n+1-2rd)$ for ultrashort $\alpha$, $|u^{-1}(\alpha)|=2d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for long $\alpha$. The outer automorphism group is cyclic of order $2$.
    Such split isotropic structure on $\mathbb{SL}_{n+1}$ corresponds to the block decomposition by $r$ strips of width $d$ from each side and, possibly, the central strip of width n+12rdn+1-2rd, both vertical and horizontal. We number these strips by integers from r-r to $r$, 0 corresponds to the central strip. The group subscheme $L$ consists of block-diagonal matrices. If $\alpha$ is short, then $U_{\alpha}$ consists of matrices differing from $1$ only in two blocks with indices $(i,j)$ and (j,i)(-j,-i) for some $0\neq|i|\neq|j|\neq 0$. If $\alpha$ is long, then $U_{\alpha}$ consists of matrices differing from $1$ only in the block with the indices (i,i)(-i,i) for some $i\neq 0$. Finally, if $\alpha$ is ultrashort, then $U_{\alpha}$ consists of matrices differing from $1$ only in the blocks with the indices $(0,i)$, (i,i)(-i,i), and (i,0)(-i,0) for some $i\neq 0$.
  • $\mathsf{B}_{n,r}$, where $1\leq r\leq n$. Here $\widetilde{\Phi}=\mathsf{B}_{n}$, $J=\{1,2,\ldots,r\}$, and $\Phi=\mathsf{B}_{r}$. The kernel of $u$ had the type $\mathsf{B}_{n-r}$, $|u^{-1}(\alpha)|=2n+1-2r$ for short $\alpha$, and $|u^{-1}(\alpha)|=1$ for long $\alpha$. The outer automorphism group is trivial.
    Such split isotropic structure on $\mathbb{SO}_{2n+1}$ is induced from the block decomposition $^{2}\!{\mathsf{A}_{2n+1,r}^{(1)}}$} on the standard representation, i.e. $\mathbb{SO}_{2n+1}\leq\mathbb{SL}_{2n+1}$ is the stabilizer of the quadratic form
    $q(\vec{x})=x_{1}x_{2n+1}+x_{2}x_{2n}+\ldots+x_{n}x_{n+1}+x_{n+1}^{2}.$
  • $\mathsf{C}_{n,r}^{(d)}$, where $d=2^{k}\mid 2n$, $rd\leq n$, $d\geq 1$, $r\geq 1$, and $n=r$ in the case $d=1$. Here $\widetilde{\Phi}=\mathsf{C}_{n}$, $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{BC}_{r}$ for $rd\leq n-1$ and $\Phi=\mathsf{C}_{r}$ for $rd=n$. The kernel of $u$ has the type $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{C}_{n-rd}$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=\frac{d(d+1)}{2}$ for long $\alpha$. The outer automorphism group is trivial.
    Such split isotropic structure on $\mathbb{Sp}_{2n}$ is induced from the block decomposition $^{2}\!{\mathsf{A}_{2n,r}^{(1)}}$} on the standard representation, i.e. $\mathbb{Sp}_{2n}\leq\mathbb{GL}_{2n}$ is the stabilizer of the symplectic form
    $B(\vec{x},\vec{y})=x_{1}y_{2n}+x_{2}y_{2n-1}+\ldots+x_{n}y_{n+1}-x_{n+1}y_{n}-x_{n+2}y_{n-1}-\ldots-x_{2n}y_{1}.$
  • $^{1}\!{\mathsf{D}_{n,r}^{(d)}}$}, where $d=2^{k}\mid 2n$, $rd\leq n$, $rd\neq n-1$, $d\geq 1$, $r\geq 1$, $n\geq 3$. Here $\widetilde{\Phi}=D_{n}$, $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{BC}_{r}$ for $rd\leq n-2$ and $d\geq 2$, $\Phi=\mathsf{B}_{r}$ for $rd\leq n-2$ and $d=1$, $\Phi=\mathsf{C}_{r}$ for $rd=n$ and $d\geq 2$, $\Phi=\mathsf{D}_{r}$ for $rd=n$ and $d=1$. The kernel of $u$ has the type $r\mathsf{A}_{d-1}+\mathsf{D}_{n-rd}$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=\frac{d(d-1)}{2}$ for long $\alpha$ (if $d=1$, then $|u^{-1}(\alpha)|=2d(n-rd)$ for short $\alpha$ and $|u^{-1}(\alpha)|=d^{2}$ for long $\alpha$). The outer automorphism group is trivial for $rd=n$ except the cases $^{1}\!{\mathsf{D}_{4,2}^{(2)}}$} and $^{1}\!{\mathsf{D}_{4,1}^{(4)}}$}; it is cyclic of order two for $^{1}\!{\mathsf{D}_{4,2}^{(2)}}$}, $^{1}\!{\mathsf{D}_{4,1}^{(4)}}$}, and for $rd\leq n-2$ except the case $^{1}\!{\mathsf{D}_{4,1}^{(2)}}$}; it is the whole triality group of order $6$ in the case $^{1}\!{\mathsf{D}_{4,1}^{(2)}}$}.
    Such split isotropic structure on $\mathbb{SO}_{2n}$ is induced from the block decomposition $^{2}\!{\mathsf{A}_{2n,r}^{(1)}}$} on the standard representation, i.e. $\mathbb{SO}_{2n}\leq\mathbb{GL}_{2n}$ is the stabilizer of the quadratic form
    $q(\vec{x})=x_{1}x_{2n+1}+x_{2}x_{2n}+\ldots+x_{n}x_{n+1}$
    and the Dickson invariant (or the determinant if $2$ is invertible).
  • $^{2}\!{\mathsf{D}_{n,r}^{(d)}}$}, where $d=2^{k}\mid 2n$, $rd\leq n-1$, $d\geq 1$, $r\geq 1$, $n\geq 3$. Here $\widetilde{\Phi}=\mathsf{D}_{n}$, $J=\{d,2d,\ldots,rd\}$ for rd<n2rd<n-2 and $J=\{d,2d,\ldots,(r-1)d,n-1,n\}$ for rd=n1rd=n-1, $\Phi=\mathsf{BC}_{r}$ for $d\geq 2$ and $\Phi=\mathsf{B}_{r}$ for $d=1$. The kernel of $u$ has the type $r\mathsf{A}_{d-1}+\mathsf{D}_{n-rd}$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=\frac{d(d-1)}{2}$ for long $\alpha$ (if $d=1$, then $|u^{-1}(\alpha)|=2d(n-rd)$ for short $\alpha$ an $|u^{-1}(\alpha)|=d^{2}$ for long $\alpha$). The outer automorphism group is cyclic of order $2$ except the case $^{2}\!{\mathsf{D}_{4,1}^{(2)}}$} and it is the triality group of order $6$ in this exceptional case.
    Such split isotropic structure on $\mathbb{SO}_{2n}$ is induced from the block decomposition $^{2}\!{\mathsf{A}_{2n,r}^{(1)}}$} on the standard representation as in the previous case, independently of whether $d=1$ or not.
Actually, all classical isotropic reductive group schemes can be described in terms of involution algebras [25].